OpenStudy (anonymous):

Factor the following polynomial: p^4 - Q^4

OpenStudy (cwrw238):

= (p^2 - Q^2)(p^2 + Q^2) = (p - Q)(p + Q)(p^2 + Q^2)

OpenStudy (anonymous):